### v1

# Uniform Norm Error Estimates for 2D Turning Point Problem

Abstract
This work presents error analysis for a finite element method applied to a two-dimensional singularly perturbed convection-diffusion turning point problem. Utilizing a layer-adapted Shishkin mesh, we prove uniform convergence in the maximum norm in the coarse layer and x-layer regions. The analysis, critically based on the properties of a discrete Green’s function, guarantees the method’s robustness and accuracy in capturing sharp solution layers. There is no work for the coerce region for two-dimensional singularly perturbed turning point problems, and also, we are getting linear order of convergence better than Stynes’ article.
Keywords: Singular Perturbation, Standard Finite Element Method, Turning Point Problem, Shishkin Mesh, Discrete Green’s Function, Pointwise Error Estimates.
1 INTRODUCTION
This paper examines a particular class of convection-diffusion boundary value problems characterized by homogeneous Dirichlet boundary conditions. The general form of the problem under consideration is
$\ -\varepsilon\Delta u+\textbf{b}(x,y).\nabla u+c(x,y)u$ $=f(x,y),\quad\text{in}\quad\Omega=(-1,1)\times(-1,1)$ (1.1a)
$u$ $=0\quad\text{on}\quad\Gamma=\partial\Omega.$ (1.1b)
Here, $\varepsilon$ is a small positive perturbation parameter $0<\varepsilon<<1$. The coefficients $b=(b_{1},b_{2})$, $c$, and the source term $f(x,y)$ are assumed to be sufficiently smooth over the closure of the domain $\bar{\Omega}=\Omega\cup\partial\Omega$.
A particular instance of this general problem (1.1) arises when it forms the focus of our study. The convection vector $b(x,y)$ is specified such that $b_{1}(x,y)=xa(x,y)$ and $b_{2}(x,y)=0\;\forall\;(x,y)\in\Omega$. This specialization leads to the following simplified problem:
$-\varepsilon\Delta u+xa(x,y)u_{x}+c(x,y)u$ $=f(x,y),\quad\text{in}\quad\Omega=(-1,1)\times(-1,1)$ (1.2a)
$u$ $=0\quad\text{on}\quad\Gamma=\partial\Omega.$ (1.2b)
$|a(x,y)|\geq\alpha>0,\quad|c(x,y)|$ $\geq 0,\quad c-\frac{1}{2}\frac{\partial b_{1}}{\partial x}\geq\gamma>0.$ (1.2c)
Here, we make a very crucial assumption about $\alpha$ and $\beta$ is that $\alpha>5/2$ and $2/3<\beta\leq 1$. Therefore, $\alpha$ cannot be arbitrarily small. In (1.2), there’s an interior layer at $x=0$. It might not have layers at the side boundaries ($x=-1\;\text{and}\;x=1$). Instead, it only has a parabolic kind of layer at the top ($y=1$) and bottom (y=1y=-1) boundaries.
This paper focuses on the finite element analysis of a two-dimensional singularly perturbed problem with a turning point along the line $x=0$, i.e., a turning line defined on the domain $\times(-1,1)$. The combination of a turning point in a two- or higher-dimensional setting presents a significant challenge for both theoretical analysis and numerical approximation. To address this, we employ a classical finite element method on a properly constructed Shishkin mesh that is adapted to the layer structures induced by the perturbation parameter and the turning point.
Researchers have been actively involved in the progression of finite difference schemes for singularly perturbed turning point problems (SPTPPs). Their work presents both single and multiple interior turning points; for example, see [3, 6, 21, 20, 12, 18, 17, 15, 13]. Berger et al. [6] and Farrell [21] conducted analyses of SPTPPs involving a single interior turning point. Their work confirmed a priori bounds on the analytic solution and its derivatives, which were crucial for demonstrating the uniform convergence of their respective numerical techniques. Specifically, Berger et al. [6] derived an $\varepsilon$-uniform error estimate for a modified El-Mistikawy-Werle finite difference scheme. In parallel, Farrell [21] achieved uniform convergence estimates for a broad category of upwinding-type schemes. In a unique contribution, Natesan and Ramanujam [15] introduced a composite numerical approach for SPTPPs that exhibit twin boundary layers. Their method combines a classical numerical scheme with an exponential fitted-difference scheme to generate an approximate solution. To implement this, the authors divided the problem domain into four distinct sub-regions and applied an asymptotic expansion approximation to solve the problem within each segment. Kadalbajoo and Gupta [20] contributed by deriving asymptotic bounds for the derivatives of the continuous solution to SPTPPs featuring twin boundary layers. These bounds were subsequently used to establish second-order $\varepsilon$-uniform convergence for a B-spline collocation method applied to a piecewise uniform layer-adapted mesh. Natesan et al. [13] proved almost first-order $\varepsilon$-uniform convergence for their proposed scheme, which was based on the classical upwind finite difference method(FDMs) implemented on a Shishkin mesh. Separately, Geng and Qian [12] introduced a stretching variable method and a reproducing kernel method for addressing SPTPPs whose solutions present twin boundary layers. In the work by Becher and Roos [3], the Richardson extrapolation method was applied to an upwind scheme, utilizing a piecewise layer-adapted mesh. This application successfully refined the order of accuracy from $N^{-2}\ln N$ to $N^{-2}(\ln N)^{2}$, where $N$ represents the number of intervals in the spatial direction.
Compared to the abundant literature on finite difference methods (FDMs), research on applying the finite element method(FEMs) to SPTPPs is limited. However, FEMs generally provide a more efficient route to higher-order convergence and require less stringent data regularity than FDMs.
Over the past decades, researchers have increasingly directed their efforts toward developing FEMs for SPTPPs. A key milestone was achieved in 1994 when Sun and Stynes [19] presented and rigorously analyzed piecewise linear Galerkin FEMs. This approximation, tested on diverse piecewise equidistant meshes, was designed for SPPs with interior turning points and solutions featuring interior layers. They successfully established parameter-uniform convergence for their method in the standard $L^{2}$ norm and a weighted energy norm. Building upon this work, Becher [23] in 2016 introduced a higher-order finite element approximation for SPPs with interior turning points. Employing Liseikin’s [25] piecewise uniform meshes, Becher’s contribution was significant in obtaining error bounds that were independent of the perturbation parameter in the energy norm. Research by Becher [4] focused on a one-dimensional SPTPPs that exhibited interior layer characteristics in its solution. The problem was discretized via a standard Galerkin FEMs, then stabilized with SDFEM on piecewise equidistant meshes, yielding parameter-independent error bounds in both the energy and SD-norms. Subsequently, in [2, 14, 9], singularly perturbed boundary turning point problems, both with and without delay, were analyzed. The application of SDFEM in these studies achieved parameter-independent convergence, approaching second order, in the maximum norm. Aasna and Rai [14] further developed the SDFEM framework for time-dependent SPPs with boundary turning points. Their strategy involved employing an $\theta$-scheme on an equidistant time mesh and SDFEM on a spatial Shishkin mesh, successfully establishing parameter-uniform stability and convergence estimates in the maximum norm for their method.
To the best of our knowledge, the pointwise error estimate of the finite element method for two-or higher-dimensional SPTPPs has received considerably less attention in the existing literature. In [7] (1.2) is considered, and a uniform energy norm estimate of order $N^{-1}\log^{2}N$ is derived for the SIPG method on a layer-adapted Shishkin mesh. In [10], Stynes used a Galerkin finite element method on a piecewise equidistant mesh to solve singular Perturbation boundary value problems of convection diffusion type in two dimensions without a turning point. That analysis indicated, uniformly convergent in the perturbation parameter of order $N^{-1}\log N$ in global energy norm and of order $N^{-1/2}\log^{3/2}N$ pointwise near the outflow boundary but not in the coarse region. Key strength is obtained by employing a discrete Green’s function, as it allows us to derive the pointwise error estimates. This observation strongly motivates our present work.
Our research seeks to investigate the pointwise error estimates of the classical finite element method on a Shishkin mesh for the problem of type (1.2). In our article, we implement a classical finite element method on a piecewise uniform mesh(known as Shishkin mesh), but we adopt some specific transition parameter along with xaxisx-axis and along with yaxisy-axis as $\lambda_{x}=min\left(\frac{2\varepsilon}{\alpha}\log\left(\frac{1}{\varepsilon}\right),\frac{1}{2}\right),\;\lambda_{y}=min\left(2\sqrt{\frac{\varepsilon}{\beta}}\log\left(\frac{1}{\varepsilon^{3/2}}\right),\frac{1}{4}\right)$ respectively to solve the convection diffusion boundary value problems with homogeneous Dirichlet boundary conditions with a turning point in the domain $(-1,1)\times(-1,1)$. We establish uniform convergence in the perturbation parameter of order $N^{-1}\log^{3}(1/\varepsilon^{3/2})$ pointwise in the coarse region and $x$-layer region.
This paper proceeds as follows: In Section 1, we establish bounds for the solution of the continuous problem (1.2). In Section 2, we present the classical finite element method, describing its key components such as the finite element space, the chosen basis functions, and the layer-adapted piecewise uniform mesh(known as the Shishkin mesh) with specific mesh transitions. Due to the symmetry along the $y$-axis only need to analyze the problem in the domain $0,1]\times[-1,1]$ along with we decompose our domain into four parts $\Omega_{c},\Omega_{f,x},\Omega_{f,y}$ and $\Omega_{f,xy}$ represents coarse region, $x$-layer region, $y$-layer region and $xy$-layer region respectively. In Section 3, we begin by recalling the discrete Green’s function, which plays a key role in deriving pointwise error estimates. We then proceed to derive $L^{2}$-norm estimates for the discrete Green’s function. A key contribution of this section is a $\varepsilon$-independent estimate for the discrete Green’s function when the source point lies in either the coarse region or the $x$-layer region. This estimate serves as a critical component in the proof of our main theorem. In Section 4, we establish $L^{\infty}$(pointwise) error estimates for the diffusion term. To achieve this, we decompose the solution into four parts: R, $F_{1}$, $F_{2}$, and $F_{12}$ corresponding to no-layer(smooth) region, $x$-layer, $y$-layer, and $xy$-layer regions, respectively. Then, we recall an interpolation theorem from [7], followed by the assumption that source points lie in coarse and $x$-layer regions, and we derived pointwise error estimates for all layer regions. In section 5, following the interpolation framework by Zarin and H.Ross, we develop $L^{\infty}$(pointwise) error estimates for the convection and reaction term. The analysis yields $\varepsilon$-independent error bounds provided the source points reside within the coarse and $x$-layer regions. (pointwise) error estimates for the convection and diffusion terms, following the interpolation framework introduced by Zarin and H. Ross. The analysis yields $\varepsilon$-independent error bounds provided the source point resides within the coarse and x-layer regions. Combining these with estimates for convection, diffusion, and reaction components, we establish the main theorem, which confirms uniform convergence in the perturbation parameter $\varepsilon$ specifically within these regions. Based on the estimates for the convection, diffusion, and reaction terms, we derive the main theorem of this work, which demonstrates uniform convergence in the perturbation parameter $\varepsilon$ of order $N^{-1}\log^{3}(1/\varepsilon^{3/2})$ pointwise specifically with these regions. In section 6, based on the theoretical, we present an example of the convection diffusion turning point problem. The double mesh principle is employed to compute the error estimates and determine the rate of convergence for various values of the perturbation parameter $\varepsilon$ and the mesh discretization parameter N. The pointwise error estimates are presented in Tables 1 and 2 when source points are within smooth and $x$-layer regions, respectively. Table 3 provides the convergence behavior in both subdomains, which is observed to be nearly first order. The numerical solution for Example 1 is depicted in the accompanying figure, thereby demonstrating the effectiveness of the proposed approach. To the best of our knowledge, there has not been any work on the pointwise error estimate for the two or higher-dimensional turning point problems.
1.1 The continuous problem
The weak formulation of the problem (1.1) is to find an unique $u\in H_{0}^{1}(\Omega)=\{u\in H^{1}(\Omega);u|_{\partial\Omega}=0\}$ such that
$\varepsilon\int_{\Omega}\nabla u.\nabla v\;dxdy+\int_{\Omega}xa(x,y)u_{x}v\;dxdy+\int_{\Omega}c(x,y)uv\;dxdy=\int_{\Omega}f(x,y)v\;dxdy\quad\forall\;v\in H_{0}^{1}(\Omega)$
$B(u,v)=(f,v)\quad\forall\;v,u\in H^{1}_{0}(\Omega),$ (1.3)
where $B(.,.)$ is defined as
$B(u,v)=(\varepsilon\nabla u,\nabla v)+(\textbf{b}.\nabla u,v)+(cu,v)\quad\forall\;u,v\in H^{1}_{0}(\Omega).$ (1.4)
The following theorem states the well-posedness of (1.3).
Theorem 1
If we assume $c(x,y)-\frac{1}{2}\frac{\partial b_{1}(x,y)}{\partial x}\geq\gamma_{1}>0\;\forall\;(x,y)\in\bar{\Omega}$ then for B(.,.) given in (1.4) the following holds
$B(u,u)\geq\frac{1}{2}\|u\|_{1,\varepsilon}^{2}\quad\forall\;u\in H^{1}_{0}(\Omega).$
Henceforth, we denote the positive function $c-\frac{1}{2}\frac{\partial b_{1}}{\partial x}$ = $c_{0}^{2}$.
Singularly perturbed differential equations are typically characterized by a small parameter $\varepsilon<<1$ multiplying some or all of the highest order terms in the differential equation. The solution of such an equation possesses interior or boundary layers or both. These are very thin regions where the derivative of the solution changes very rapidly. Outside these layers, the solution smoothly approximates that of the lower-order reduced problem (obtained by setting $\varepsilon=0$ in (1.2)). In some situations, such as nonlinearities or discontinuous source terms or a sign change of the convection coefficient, interior layers can also be present within the domain. As $\varepsilon\to 0$, the solution’s derivatives within these layers often scale inversely with powers of $\varepsilon$, indicating a significant loss of uniform regularity across the domain. For the problem (1.1), turning points in SPPs can be identified by a sign change in the convective coefficient within the domain. At a turning point, the convective term locally vanishes or becomes very small. This means that diffusion, even if small (due to $\varepsilon$), can locally become relatively more dominant, influencing the solution’s behavior. Both from a theoretical and computational perspective, turning point SPPs pose a greater challenge than standard SPPs. Standard numerical methods designed for convection-dominated problems may struggle to accurately resolve both the boundary and interior layers, often leading to spurious oscillations or significant errors if not specifically adapted (e.g., using fitted meshes).
To proceed with our analysis, precise bounds on specific derivatives of the solution $u(x,y)$ to the problem (1.2) are indispensable. This is discussed in the following theorem.
Theorem 2 [1] Let $f\in C^{3,\beta}(\bar{\Omega})$ where $\beta\in(2/3,1]$ and $f$ satisfy the compatibility condition f(1,1)=f(1,1)=f(1,1)=f(1,1)=0f(-1,-1)=f(1,1)=f(1,-1)=f(-1,1)=0, then the exact solution of given problem can be broken down as $u=R+F_{1}+F_{2}+F_{12}$, where
$R=$ $\text{Regular part of the solution},$
[…]
$F_{12}=$ Part of the solution containing both interior and boundary layers near
$\text{the points (0,1) and (0,-1)}.$
where for all $(x,y)\in\bar{\Omega}$ and $0\leq i+j\leq 2$, we have
$\left|\frac{\partial^{i+j}R(x,y)}{\partial x^{i}\partial y^{j}}\right|$ $\leq C,$
$\left|\frac{\partial^{i+j}F_{1}(x,y)}{\partial x^{i}\partial y^{j}}\right|$ $\leq C\varepsilon^{-i}e^{-\frac{\alpha x}{\varepsilon}},$
$\left|\frac{\partial^{i+j}F_{2}(x,y)}{\partial x^{i}\partial y^{j}}\right|$ $\leq C\varepsilon^{-j/2}\left(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\right),$
$\left|\frac{\partial^{i+j}F_{12}(x,y)}{\partial x^{i}\partial y^{j}}\right|$ $\leq C\varepsilon^{-i-j/2}e^{-\frac{\alpha x}{\varepsilon}}\left(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\right).$
2 A classical discretization on a Shishkin mesh
We denote the $L^{2}(\Omega)$ norm by $\|.\|$ and the energy norm by
$\|v\|^{2}_{1,\varepsilon}=\varepsilon\|\nabla v\|^{2}+\|v\|^{2}\quad\forall\;v\in H^{1}(\Omega)$ (2.1)
where $H^{1}(\Omega)$ is the usual Sobolev space of functions whose first-order partial derivatives lie in $L^{2}(\Omega)$. We discretize (1.3) by means of the standard finite element method on a special rectangular mesh. Let $\mathcal{T}_{h}$${\{\tau_{i}^{h}\}}_{i=1}^{N}$ be a family of edge-to-edge rectangle’s of $\Omega$ with the parameter $h$.
$S_{h}=\{\mathcal{X}\in\mathcal{C}^{0}(\Omega),\mathcal{X}=0\;\text{on}\;\partial\Omega,\mathcal{X}|_{\tau_{i}^{h}}\;\text{linear in x and y}.\}$
The standard finite element discretization of (1.3) is:
Find $u_{h}\in S_{h}$ such that
$B_{h}(u_{h},v_{h})=(f,v_{h})\quad\forall\;v_{h}\in S_{h},$ (2.2)
where $B_{h}(.,.)$ is the bilinear form defined as
$B_{h}(u_{h},v_{h})=(\varepsilon\nabla u_{h},\nabla v_{h})+(\textbf{b}.\nabla u_{h},v_{h})+(cu_{h},v_{h})\quad\forall\;u_{h},v_{h}\in S_{h}.$ (2.3)
To effectively address the layering phenomenon inherent in the problem’s solution, we employ a layer-adapted mesh of Shishkin type, constructed independently along both the $x$ and $y$ spatial directions. For an even positive integer $N\geq 4$, the crucial mesh transition parameters in the $x$ and $y$ directions are denoted by $\lambda_{x}$ and $\lambda_{y}$, respectively. Note that these newly defined transition parameters are one of the novel contributions of this article. They are defined as,
$\lambda_{x}=min\left(\frac{2\varepsilon}{\alpha}\log\left(\frac{1}{\varepsilon}\right),\frac{1}{2}\right),\quad\lambda_{y}=min\left(2\sqrt{\frac{\varepsilon}{\beta}}\log\left(\frac{1}{\varepsilon^{3/2}}\right),\frac{1}{4}\right),$
Mesh points along the $x$-axis and $y$-axis are,
$x_{j}=\begin{cases}\frac{2\lambda_{x}}{N}i,&i=0,1,\dots,N/2,\\ \lambda_{x}+\frac{2}{N}(1-\lambda_{x})(i-\frac{N}{2}),&i=N/2+1,\dots,N.\end{cases}$ (2.4)
and
$y_{j}=\begin{cases}-1+\frac{4j}{N}\lambda_{y},&j=0,1,\dots,\frac{N}{4},\\ -1+\lambda_{y}+\frac{2}{N}(2-2\lambda_{y})(j-\frac{N}{4}),&j=\frac{N}{4}+1,\dots,\frac{3N}{4},\\ 1-4\lambda_{y}(1-\frac{j}{N}),&j=\frac{3N}{4}+1,\dots,N.\end{cases}$ (2.5)
Figure 1: Shishkin mesh
Figure 2: Domain $\Omega$ and subdomains
In the $x$-direction, the interval $[0,1]$ is bisected into two sub-intervals $0,\lambda_{x}]$ and $\lambda_{x},1]$. Each of these sub-intervals is further partitioned such that its closure contains $N/2+1$ points, denoted as $X_{K}^{N}$. A symmetric partitioning is implicitly applied to [1,0][-1,0] to cover the full $x$ range. For the $y$-direction, the interval [1,1][-1,1] is divided into three sub-intervals: $1,-1+\lambda_{y}]$, $1+\lambda_{y},1-\lambda_{y}]$ and $1-\lambda_{y},1]$. These sub-intervals are then uniformly discretized. Specifically, their closures are partitioned into $N/4+1$, $N/2+1$, and $N/4+1$ points, respectively, forming the set $Y_{J}^{N}$. The global mesh points $\Omega^{N}$ are defined as the tensor product of these one-dimensional partitions:
$\Omega^{N}=\{(x_{i},y_{j})\in\Omega:x_{i}\in X_{K}^{N},y_{j}\in Y_{J}^{N},i=0,1,\dots,N,j=0,1,\dots,N\},\quad N\in\mathbb{N}.$
Axiparallel lines passing through the mesh points generate a collection of rectangular mesh elements, denoted as $\mathcal{J}_{N}$. This subdivision implicitly creates different regions within $\Omega$, often categorized as coarse ($\Omega_{c}$) and fine ($\Omega_{f,x}\;$, $\Omega_{f,y}$, $\Omega_{f,xy}$) mesh region. Note that in the coarse region, the solution is expected to behave nicely, and in the layer regions, the solution is expected to exhibit layers. The mesh widths in the layer regions and coarse region are calculated as:
$h_{x}$ $=2\lambda_{x}N^{-1},\quad\quad H_{x}=2(1-\lambda_{x})N^{-1}$ (2.6)
$h_{y}$ $=4\lambda_{y}N^{-1},\quad\quad H_{y}=4(1-\lambda_{y})N^{-1}$ (2.7)
Here, $h_{x}$, $h_{y}$ denote the mesh widths within the layer regions in the $x$ and $y$ directions, respectively, while $H_{x}$, $H_{y}$ represent the mesh widths outside these regions. Due to the symmetry of the problem with respect to the $y$-axis, we analyze our problem in the reduced domain $\Omega=[0,1]\times[-1,1]$[7]. Within this domain, we perform a decomposition into four non-overlapping subregions, tailored to distinguish between areas of expected smooth behavior and those containing layers. This partition is defined as:
$\Omega_{c}=[\lambda_{x},1]\times[-1+\lambda_{y},1-\lambda_{y}],\quad\Omega_{f,x}=[0,\lambda_{x}]\times[-1+\lambda_{y},1-\lambda_{y}]$
$\Omega_{f,y}=([-1,-1+\lambda_{y}]\cup[1-\lambda_{y},1])\times[\lambda_{x},1],\quad\Omega_{f,xy}=[0,\lambda_{x}]\times([-1,-1+\lambda_{y}]\cup[1-\lambda_{y},1])$
3 Estimates for Discrete Green’s function
This section is devoted to finding out some important estimates for the discrete Green’s function corresponding to the differential operator in (1.1). These estimates play a fundamental role in deriving the pointwise error estimates. We derive the estimates of the discrete Green’s function in $L^{2}$ and energy norms. We note here that the Green’s function, or the discrete Green’s function, contains information about the source point. Therefore, these estimates play a crucial role in determining whether we can obtain an $\varepsilon$-independent estimate or not. We shall see that if the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then we obtain an $\varepsilon$-independent estimate for the Green’s function in energy and $L^{2}$ norms, and we can obtain a parameter-uniform $L^{\infty}$ norm error estimate.
Lemma 3.1 When the source point $(x_{m},y_{n})\in\Omega_{c}$, then $\|G_{mn}\|_{1,\varepsilon}\leq CN.$
Proof. From the coercivity of the bilinear form we have,
$C_{1}\|G_{mn}\|^{2}_{1,\varepsilon}$ $\leq B(G_{mn},G_{mn})=G_{mn}(x_{m},y_{n})\leq|G_{mn}(x_{m},y_{n})|$
$\leq CN^{2}\|G_{mn}\|_{L^{1}(R_{mn})}\quad\quad\quad(\text{by inverse estimate})$
Where $R_{mn}$ is the union of those rectangles whose common vertex is $(x_{m},y_{n})$. Therefore
$C_{1}\|G_{mn}\|^{2}_{1,\varepsilon}$ $\leq CN^{2}\|G_{mn}\|_{L^{1}(R_{mn})}$
$\leq CN^{2}|R_{mn}|^{1/2}\|G_{mn}\|_{L^{2}({R_{mn}})}$
$\leq CN^{2}\frac{1}{N}\|G_{mn}\|\leq CN\|G_{mn}\|_{{1,\varepsilon}}$
$\|G_{mn}\|_{{1,\varepsilon}}$ $\leq CN.$
Corollary 1 When $(x_{m},y_{n})\in\Omega_{c}$ then $\|G_{mn}\|\leq CN.$
Lemma 3.2 When $(x_{m},y_{n})\in\Omega_{f,x}$ then $\|G_{mn}\|_{1,\varepsilon}\leq CN^{1/2}\log^{1/2}(1/\varepsilon).$
Proof. From the coercivity of the bilinear form we have,
$C_{1}\|G_{mn}\|^{2}_{1,\varepsilon}\leq B(G_{mn},G_{mn})$ $=G_{mn}(x_{m},y_{n})=-\int_{t=x_{m}}^{1}G_{mn,x}(t,y_{n})\;dt$
$\leq\int_{x_{m}}^{1}|G_{mn,x}(t,y_{n})|\;dt$
$\leq CN\int_{y_{n-1}}^{y_{n}}\int_{x_{m}}^{1}|G_{mn,x}(t,s)|\;dsdt$
$\leq CN\frac{\varepsilon^{1/2}\log^{1/2}(1/\varepsilon)}{N^{1/2}}\|G_{mn,x}\|$
$\leq CN^{1/2}\log^{1/2}(1/\varepsilon)\|G_{mn}\|_{1,\varepsilon}.$
$\|G_{mn}\|_{1,\varepsilon}$ $\leq CN^{1/2}\log^{1/2}(1/\varepsilon).$
Corollary 2 When $(x_{m},y_{n})\in\Omega_{f,x}$ then $\|G_{mn}\|\leq C\varepsilon^{-1/2}\log^{-1/2}(1/\varepsilon).$
Remark 1 Due to the $\varepsilon$-dependent estimate for $\|G_{mn}\|$ when $(x_{m},y_{n})\in\Omega_{f,y}\cup\Omega_{f,xy}$ it is not possible to obtain $\varepsilon$-independent error estimate. Note that when $(x_{m},y_{n})\in\Omega_{f,y}$ we are able to obtain $\|G_{mn}\|\leq C\varepsilon^{-1/4}\log^{-1/2}(1/\varepsilon)$ and when $(x_{m},y_{n})\in\Omega_{f,xy}$ then $\|G_{mn}\|\leq C\varepsilon^{-1/4}$. We are not able to obtain a parameter uniform estimate for $\|G_{mn}\|_{1,\varepsilon}$ when the source point $(x_{m},y_{n})\in\Omega_{f,y}\cup\Omega_{f,xy}$, which implies the $\varepsilon$-dependent error estimate in $L^{\infty}$ norm, which is also reflected in the numerical examples.
Theorem 2
[7] Assume that $K\in\mathcal{J}_{N}$ is a mesh rectangle with sides that are perpendicular to the coordinate axes. Assume that $w\in H^{3}(K)$ and that $w^{I}$ is its nodal bilinear interpolate on $K$. Hence, we obtain for any bilinear function $v^{N}$ defined on $K$
$\left|\int_{K}(Iu-u)_{x}v_{x}^{N}\,dxdy\right|$ $\leq Ch_{y,K}^{2}\|u_{xyy}\|_{L^{2}(K)}\|v_{x}^{N}\|_{L^{2}(K)},$
$\left|\int_{K}(Iu-u)_{y}v_{y}^{N}\,dxdy\right|$ $\leq Ch_{x,K}^{2}\|u_{xxy}\|_{L^{2}(K)}\|v_{y}^{N}\|_{L^{2}(K)}.$
Lemma 4.1 If the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then $\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega}\leq\frac{C(\log(1/\varepsilon^{3/2}))^{3}}{N}.$
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$.
$\varepsilon(\nabla(R-IR),\nabla G_{mn})$ $=\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{f,xy}}+\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{f,x}}$
$+\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{f,y}}+\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{c}}$ (3.1)
Now, consider the first term of the R.H.S. of (3.1) in the $x$-direction $\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,xy}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{R}_{xyy}\|_{\Omega_{f,xy}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,xy}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{|\Omega_{f,xy}|}\>\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\frac{\varepsilon^{9/4}(\log(1/\varepsilon^{3/2}))^{3}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$ (3.2)
Now, consider the first term of the R.H.S. of (3.1) in the $y$-direction $\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,xy}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,xy}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,xy}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\sqrt{|\Omega_{f,xy}|}\;\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\frac{\varepsilon^{13/4}(\log(1/\varepsilon^{3/2}))^{3}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$ (3.3)
Now, consider the second term of the R.H.S. of (3.1) in the $x$-direction $\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,x}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{R}_{xyy}\|_{\Omega_{f,x}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,x}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{1}{N}\right)^{2}\sqrt{|\Omega_{f,x}|}\>\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\left(\frac{1}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\;N$
$\leq C\frac{\varepsilon(\log(1/\varepsilon))^{1/2}}{N}\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N}.$ (3.4)
Now, consider the second term of the R.H.S. of (3.1) in the $y$-direction $\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,x}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,x}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,x}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\sqrt{|\Omega_{f,x}|}\;\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\;N$
$\leq C\frac{\varepsilon^{3}(\log(1/\varepsilon))^{5/2}}{N}\leq C\frac{(\log(1/\varepsilon))^{5/2}}{N}.$ (3.5)
Now, consider the third term of the R.H.S. of (3.1) in the $x$-direction $\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,y}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{R}_{xyy}\|_{\Omega_{f,y}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,y}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{|\Omega_{f,y}|}\>\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\frac{\varepsilon^{7/4}(\log(1/\varepsilon^{3/2}))^{5/2}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}.$ (3.6)
Now, consider the third term of the R.H.S. of (3.1) in the $y$-direction $\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,y}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,y}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,y}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{1}{N}\right)^{2}\sqrt{|\Omega_{f,y}|}\;\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\left(\frac{1}{N}\right)^{2}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\frac{\varepsilon^{3/4}(\log(1/\varepsilon^{3/2}))^{1/2}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{1/2}}{N}.$ (3.7)
Now, consider the fourth term of the R.H.S. of (3.1) $\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{c}}$. We obtain,
$\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{c}}$ $\leq\varepsilon^{1/2}\|\nabla(R-IR)\|\varepsilon^{1/2}\|G_{mn,x}\|_{\Omega_{c}}$
$\leq\varepsilon^{1/2}CN^{-2}N$
$\leq C\varepsilon^{1/2}\frac{1}{N}\leq\frac{C}{N}.$ (3.8)
Applying similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega}\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, we have
$\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega}\leq\frac{C(\log(1/\varepsilon^{3/2}))^{3}}{N}$
Lemma 4.2 When $\alpha\geq 3/2$ and the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}$.
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$.
$\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})$ $=\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega_{f,xy}}+\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega_{f,x}}$
$+\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega_{f,y}}+\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega_{c}}$ (3.9)
Now, consider the first term of the R.H.S. of (3.9) in the $x$-direction $\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,xy}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{F_{1}}_{xyy}\|_{\Omega_{f,xy}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,xy}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{|\Omega_{f,xy}|}\frac{1}{\varepsilon}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,xy}})}\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\left(\frac{\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$ (3.10)
Now, consider the first term of the R.H.S. of (3.9) in the $y$-direction $\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,xy}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,xy}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,xy}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\sqrt{|\Omega_{f,xy}|}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,xy}})}\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\left(\frac{\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$ (3.11)
Now, consider the second term of the R.H.S. of (3.9) in the $x$-direction $\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,x}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{F_{1}}_{xyy}\|_{\Omega_{f,x}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,x}}}$
$\leq C\sqrt{\varepsilon}\frac{1}{N^{2}}\sqrt{|\Omega_{f,x}|}\frac{1}{\varepsilon}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,x}})}\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\frac{1}{\varepsilon N^{2}}\sqrt{\varepsilon\log(1/\varepsilon)}\;N\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N}.$ (3.12)
Now, consider the second term of the R.H.S. of (3.9) in the $y$-direction $\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,x}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,x}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,x}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\sqrt{|\Omega_{f,x}|}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,x}})}\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\left(\frac{\log(1/\varepsilon)}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\;N$
$\leq C\varepsilon^{1/2}\frac{(\log(1/\varepsilon))^{3}}{N}\leq C\frac{(\log(1/\varepsilon))^{3}}{N}.$ (3.13)
Now, consider the third term of the R.H.S. of (3.9) in the $x$-direction $\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,y}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{F_{1}}_{xyy}\|_{\Omega_{f,y}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,y}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{|\Omega_{f,y}|}\frac{1}{\varepsilon}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,y}})}\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\frac{(\log(1/\varepsilon^{3/2}))^{2}}{N^{2}}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\varepsilon^{3/4}\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}.$ (3.14)
Now, consider the third term of the R.H.S. of (3.9) in the $y$-direction $\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,y}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,y}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,y}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{1}{N}\right)^{2}\sqrt{|\Omega_{f,y}|}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,y}})}\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\frac{1}{N^{2}}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,y}})}\;N$
$\leq C\varepsilon^{-3/2}\frac{\log(1/\varepsilon^{3/2})}{N}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,y}})}.$ (3.15)
Note that when $(x,y)\in\Omega_{f,y}$, then $x\in[\varepsilon\log(1/\varepsilon),1]$ and $y\in[-1,-1+\lambda_{y}]\cup[1-\lambda_{y},1]$.
If $\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}=e^{-\frac{\alpha x}{\varepsilon}}|_{x=\lambda_{x}}$
$\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}$ $=e^{-\frac{\alpha}{\varepsilon}\varepsilon\log\left(\frac{1}{\varepsilon}\right)}=e^{-\alpha\log\left(\frac{1}{\varepsilon}\right)}=e^{\log\left(\frac{1}{\varepsilon^{-\alpha}}\right)}=\varepsilon^{\alpha}.$ (3.16)
If $\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}=e^{-\frac{\alpha x}{\varepsilon}}|_{x\neq\lambda_{x}}$ and $x_{0}\in(\lambda_{x},1]$, we obtain
$\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}$ $=e^{-\frac{\alpha x_{0}}{\varepsilon}}=\frac{1}{e^{\frac{\alpha x_{0}}{\varepsilon}}}=\frac{1}{1+\frac{\alpha x_{0}}{1!\varepsilon}+\frac{\alpha^{2}x_{0}^{2}}{2!\varepsilon^{2}}+\frac{\alpha^{3}x_{0}^{3}}{3!\varepsilon^{3}}+\dots}\leq n!\varepsilon^{n}$
$\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}$ $\leq C\varepsilon^{\alpha}.$ (3.17)
Combining (3.15), (3.16) and (3.17), we obtain
$\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,y}}$ $\leq C\varepsilon^{-3/2}\frac{\log(1/\varepsilon^{3/2})}{N}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,y}})}$
$\leq C\varepsilon^{-3/2}\frac{\log(1/\varepsilon^{3/2})}{N}\varepsilon^{\alpha}\leq C\frac{\log(1/\varepsilon^{3/2})}{N}\quad\text{when}\;\alpha\geq 3/2.$ (3.18)
Now, consider the fourth term of the R.H.S. of in the $x$-direction (3.9) $\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{c}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{c}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{F_{1}}_{xyy}\|_{\Omega_{c}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{c}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{1}{N}\right)^{2}\sqrt{|\Omega_{c}|}\frac{1}{\varepsilon}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{c}})}\|G_{mn}\|_{1,\varepsilon}$
$\leq C\frac{1}{N^{2}}\frac{1}{\sqrt{\varepsilon}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{c}})}N\leq C\varepsilon^{-1/2}\frac{1}{N}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{c}})}.$ (3.19)
Note that when $(x,y)\in\Omega_{c}$, then $x\in[\varepsilon\log(1/\varepsilon),1]$ and $y\in[-1+\lambda_{y},1-\lambda_{y}]$.
when $x\in[\lambda_{x},1]$
$x$ $\geq\lambda_{x}\implies\frac{-\alpha x}{\varepsilon}\leq\frac{-\alpha\lambda_{x}}{\varepsilon}$
$e^{\frac{-\alpha x}{\varepsilon}}$ $\leq e^{\frac{-\alpha\lambda_{x}}{\varepsilon}}=e^{-\frac{\alpha}{\varepsilon}\varepsilon\log\left(\frac{1}{\varepsilon}\right)}=e^{-\alpha\log\left(\frac{1}{\varepsilon}\right)}=e^{\log\left(\frac{1}{\varepsilon^{-\alpha}}\right)}=\varepsilon^{\alpha}$
$\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}$ $\leq C\varepsilon^{\alpha}.$ (3.20)
Combining (3.19) and (3.20), we obtain
$\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{c}}$ $\leq C\varepsilon^{-1/2}\frac{1}{N}\leq\varepsilon^{\alpha-1/2}\frac{1}{N}\leq\frac{C}{N}.\quad\text{when}\;\alpha\geq 1/2.$ (3.21)
Now, consider the fourth term of the R.H.S. of in the $y$-direction (3.9) $\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{c}}$. We obtain,
$\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{c}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{c}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{c}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{1}{N}\right)^{2}\sqrt{|\Omega_{c}|}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{c}})}\|G_{mn}\|_{1,\varepsilon}$
$\leq C\frac{1}{N^{2}}\frac{1}{\varepsilon^{3/2}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{c}})}N\leq C\varepsilon^{-3/2}\frac{1}{N}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{c}})}.$ (3.22)
Combining (3.22) and (3.20), we obtain
$\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{c}}$ $\leq C\frac{\varepsilon^{\alpha-3/2}}{N}\leq\frac{C}{N}.\quad\text{when}\;\alpha\geq 3/2.$ (3.23)
Applying similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, we have
$\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
Lemma 4.3 When $\beta\geq 1/3$ and the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}$.
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$.
$\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})$ $=\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega_{f,xy}}+\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega_{f,x}}$
$+\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega_{f,y}}+\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega_{c}}$ (3.24)
Now, consider the first term of the R.H.S. of (3.24) in the $x$-direction $\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,xy}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{x}\|_{\Omega_{f,xy}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,xy}}$
$\leq C\varepsilon^{1/2}\;h_{y}^{2}\|{F_{2}}_{xyy}\|_{\Omega_{f,xy}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\varepsilon^{1/2}\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\frac{1}{\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,xy})}$
$\quad\sqrt{|\Omega_{f,xy}|}\;N\leq C\varepsilon^{1/2}\frac{(\log(1/\varepsilon^{3/2})^{2}}{N}\varepsilon^{3/4}\log(1/\varepsilon^{3/2})N$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$ (3.25)
Now, consider the first term of the R.H.S. of (3.24) in the $y$- direction $\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,xy}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{y}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,xy}}$
$\leq C\varepsilon^{1/2}\;h_{x}^{2}\|{F_{2}}_{yxx}\|_{\Omega_{f,xy}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{1/2}}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,y})}$
$\quad\sqrt{|\Omega_{f,xy}|}\;N\leq C\frac{1}{N}\varepsilon^{3/4}\log(1/\varepsilon^{3/2})$
$\leq C\varepsilon^{3/4}\frac{\log(1/\varepsilon^{3/2})}{N}\leq C\frac{\log(1/\varepsilon^{3/2})}{N}.$ (3.26)
Now, consider the second term of the R.H.S. of (3.24) in the $x$-direction $\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,x}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{x}\|_{\Omega_{f,x}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,x}}$
$\leq C\varepsilon^{1/2}\;h_{y}^{2}\|{F_{2}}_{xyy}\|_{\Omega_{f,x}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}\sqrt{|\Omega_{f,x}|}\;N$
$\leq C\varepsilon^{1/2}\frac{1}{N\varepsilon}\varepsilon^{1/2}(\log(1/\varepsilon)^{1/2}\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N}.$ (3.27)
Now, consider the second term of the R.H.S. of (3.24) in the $y$-direction $\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,x}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{y}\|_{\Omega_{f,x}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,x}}$
$\leq C\varepsilon^{1/2}\;h_{x}^{2}\|{F_{2}}_{yxx}\|_{\Omega_{f,x}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\varepsilon^{1/2}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\frac{1}{\varepsilon^{1/2}}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}$
$\quad\sqrt{|\Omega_{f,x}|}\;N\leq C\frac{1}{N}\varepsilon^{2}(\log(1/\varepsilon))^{2}\leq C\frac{(\log(1/\varepsilon))^{2}}{N}.$ (3.28)
Now, consider the third term of the R.H.S. of (3.24) in the $x$-direction $\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{f,y}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{x}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,y}}$
$\leq C\varepsilon^{1/2}\;h_{y}^{2}\|{F_{2}}_{xyy}\|_{\Omega_{f,y}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\varepsilon^{1/2}\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\frac{1}{\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,y})}$
$\quad\sqrt{|\Omega_{f,y}|}\;N\leq C\varepsilon^{1/2}\frac{(\log(1/\varepsilon^{3/2})^{2}}{N}\varepsilon^{1/4}(\log(1/\varepsilon^{3/2})^{1/2}$
$\leq C\varepsilon^{3/4}\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}.$ (3.29)
Now, consider the third term of the R.H.S. of (3.24) $\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{f,y}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{y}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,y}}$
$\leq C\varepsilon^{1/2}\;h_{x}^{2}\|{F_{2}}_{yxx}\|_{\Omega_{f,y}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{1/2}}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,y})}$
$\quad\sqrt{|\Omega_{f,y}|}\;N\leq C\frac{1}{N}\varepsilon^{1/4}(\log(1/\varepsilon^{3/2})^{1/2}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{1/2}}{N}.$ (3.30)
Now, consider the fourth term of the R.H.S. of (3.24) in the $x$-direction $\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{c}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{c}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{x}\|_{\Omega_{c}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{c}}$
$\leq C\varepsilon^{1/2}\;h_{y}^{2}\|{F_{2}}_{xyy}\|_{\Omega_{c}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\frac{\sqrt{\varepsilon}}{\varepsilon N^{2}}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;N$
$\leq C\frac{1}{\varepsilon^{1/2}N}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{c})}.$ (3.31)
when $(x,y)\in\Omega_{c}$, then $x\in[\varepsilon\log\left(\frac{1}{\varepsilon}\right),1]$ and $y\in\left[-1+\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right),1-\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)\right]$.
Then say $y=-1+\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})$, we obtain
$1-y$ $=1-\left(-1+\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)\right)=2-\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)$
$1-y$ $\leq 2.\quad\quad\text{and }\quad 1+y\geq 1.$ (3.32)
For sufficiently small $\varepsilon>0$, we obtain
$e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}$ $\leq e^{-\frac{\beta}{\sqrt{\varepsilon}}}=\frac{1}{e^{\frac{\beta}{\sqrt{\varepsilon}}}}=\frac{1}{1+\frac{\beta}{1!\sqrt{\varepsilon}}+\frac{\beta^{2}}{2!\sqrt{\varepsilon}^{2}}+\frac{\beta^{3}}{3!\sqrt{\varepsilon}^{3}}+\dots}$
$e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}$ $\leq C\varepsilon^{\eta}\quad\quad\text{for any}\quad\eta>0.$ (3.33)
Also $1+y=\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})$, we obtain
$e^{-\frac{\beta}{\sqrt{\varepsilon}}\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)}$ $\leq e^{-\beta\log\left(\frac{1}{\varepsilon^{3/2}}\right)}\leq e^{\log\left(\frac{1}{\varepsilon^{\frac{-3\beta}{2}}}\right)}\leq e^{\log\left(\varepsilon^{\frac{3\beta}{2}}\right)}\leq\varepsilon^{\frac{3\beta}{2}}.$ (3.34)
Now, combining (3.31),(3.33) and (3.34), we obtain
$\varepsilon((F_{2}-IF_{2})_{x},G_{mn,x})_{\Omega_{c}}$ $\leq C\frac{1}{\varepsilon^{1/2}N}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{c})}$
$\leq C\frac{\varepsilon^{-1/2}}{N}\varepsilon^{\frac{3\beta}{2}}\leq\frac{C}{N}.\quad\text{when}\;\beta\geq 1/3.$ (3.35)
Now, consider the fourth term of the R.H.S. of (3.24) in the $y$-direction $\varepsilon((F_{2}-IF_{2})_{y},G_{mn,x})_{\Omega_{c}}$. We obtain,
$\varepsilon((F_{2}-IF_{2})_{y},G_{mn,y})_{\Omega_{c}}$ $\leq\sqrt{\varepsilon}\|(F_{2}-IF_{2})_{y}\|_{\Omega_{c}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{c}}$
$\leq C\varepsilon^{1/2}\;h_{x}^{2}\|{F_{2}}_{yxx}\|_{\Omega_{c}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{1/2}}N\leq\frac{C}{N}.$ (3.36)
Applying similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})\leq C\frac{(\log(1/\varepsilon^{3/2})^{7/2}}{N^{3/2}}.$
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, we have
$\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
Lemma 4.4 When $\alpha>2$, $\beta>2/3$ and the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}$.
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$.
$\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})$ $=\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega_{f,xy}}+\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega_{f,x}}$
$+\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega_{f,y}}+\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega_{c}}$ (3.37)
Now, consider the first term of the R.H.S. of in the $x$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,xy}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{x}\|_{\Omega_{f,xy}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,xy}}$
$\leq C\varepsilon^{1/2}\;h_{y}^{2}\|{F_{12}}_{xyy}\|_{\Omega_{f,xy}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,xy})}$
$\quad\frac{\varepsilon^{1/2}}{\varepsilon^{2}}\sqrt{|\Omega_{f,xy}|}\;N\leq C\varepsilon^{1/2}\frac{(\log(1/\varepsilon^{3/2})^{2}}{\varepsilon N}\varepsilon^{3/4}\log(1/\varepsilon^{3/2})$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N}.$ (3.38)
Now, consider the first term of the R.H.S. of in the $y$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,xy}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{y}\|_{\Omega_{f,xy}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,xy}}$
$\leq C\varepsilon^{1/2}\;h_{x}^{2}\|{F_{12}}_{yxx}\|_{\Omega_{f,xy}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\varepsilon^{1/2}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\frac{\varepsilon^{3/4}}{\varepsilon^{2}}\log(1/\varepsilon^{3/2})N$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$ (3.39)
Now, consider second term of the R.H.S. of in the $x$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,x}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{x}\|_{\Omega_{f,x}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,x}}$
$\leq C\varepsilon^{1/2}\;h_{y}^{2}\|{F_{12}}_{xyy}\|_{\Omega_{f,x}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,x})}$
$\quad\sqrt{|\Omega_{f,x}|}\;N\leq C\frac{(\varepsilon\log(1/\varepsilon))^{1/2}}{\varepsilon^{3/2}N}$
$\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,x})}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}.$ (3.40)
$\text{Note that}\quad\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,x})}=1.$ (3.41)
Consider $\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}.$
Note that when $(x,y)\in\Omega_{f,x}$, then $x\in[0,\varepsilon\log(1/\varepsilon)]$ and $y\in[-1+\lambda_{y},1-\lambda_{y}]$.
If $\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}=\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}$.
Let $\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}=e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y_{0})}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y_{0})}$ where $y_{0}\in[-1+\lambda_{y},1-\lambda_{y}]$.
Let $y_{0}=-1+\lambda_{y}$ then $1+y_{0}=\lambda_{y}=\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)$
$e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y_{0})}$ $=e^{-\frac{\beta}{\sqrt{\varepsilon}}\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)}=e^{-\beta\log\left(\frac{1}{\varepsilon^{3/2}}\right)}=e^{\log\left(\frac{1}{\varepsilon^{-\frac{3\beta}{2}}}\right)}=\varepsilon^{\frac{3\beta}{2}}.$ (3.42)
and $1-y_{0}=1-(-1+\lambda_{y})=2-\lambda_{y}=2-\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)\geq 1$ for sufficiently small $\varepsilon>0$.
$e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y_{0})}$ $=\frac{1}{e^{\frac{\beta}{\sqrt{\varepsilon}}(1-y_{0})}}=\frac{1}{1+\frac{\beta}{1!\sqrt{\varepsilon}}+\frac{\beta^{2}}{2!\sqrt{\varepsilon}^{2}}+\frac{\beta^{3}}{3!\sqrt{\varepsilon}^{3}}+\dots}\leq C\varepsilon^{\frac{3}{2}\beta}.$ (3.43)
Combining (3.40), (3.41), (3.42) and (3.43), we obtain
$\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,x}}$ $\leq C\frac{(\log(1/\varepsilon))^{1/2}}{\varepsilon N}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,x})}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}$
$\leq C\frac{(\log(1/\varepsilon))^{1/2}}{\varepsilon N}\varepsilon^{\frac{3}{2}\beta}\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N}.\quad\text{when}\;\beta>2/3.$ (3.44)
Now, consider second term of the R.H.S. of in the $y$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,x}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{y}\|_{\Omega_{f,x}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,x}}$
$\leq C\varepsilon^{1/2}\;h_{x}^{2}\|{F_{12}}_{yxx}\|_{\Omega_{f,x}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\frac{\varepsilon^{1/2}}{\varepsilon^{5/2}}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,x})}$
$\quad\sqrt{|\Omega_{f,x}|}\;N\leq C\frac{1}{N}\varepsilon^{1/2}(\log(1/\varepsilon))^{5/2}\leq C\frac{(\log(1/\varepsilon))^{5/2}}{N}.$ (3.45)
Now, consider the third term of the R.H.S. of in the $x$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,y}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{x}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,y}}$
$\leq C\varepsilon^{1/2}\;h_{y}^{2}\|{F_{12}}_{xyy}\|_{\Omega_{f,y}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,y})}$
$\quad\frac{\varepsilon^{1/2}}{\varepsilon^{2}}\sqrt{|\Omega_{f,y}|}\;N\leq C\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{\varepsilon^{1/4}N}e^{\alpha}$
$\leq C\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}\quad\text{when}\;\alpha>1/4.$ (3.46)
Now, consider the third term of the R.H.S. of in the $y$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,y}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{y}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,y}}$
$\leq C\varepsilon^{1/2}\;h_{x}^{2}\|{F_{12}}_{yxx}\|_{\Omega_{f,y}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{5/2}}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,y})}\sqrt{|\Omega_{f,y}|}\;N$
$\leq C\frac{(\log(1/\varepsilon^{3/2}))^{1/2}}{\varepsilon^{7/4}N}e^{\alpha}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{1/2}}{N}\quad\text{when}\;\alpha>7/4.$ (3.47)
Now, consider the fourth term of the R.H.S. of in the $x$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{c}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{c}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{x}\|_{\Omega_{c}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{c}}$
$\leq C\varepsilon^{1/2}\;h_{y}^{2}\|{F_{12}}_{xyy}\|_{\Omega_{c}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;N$
$\leq C\frac{1}{\varepsilon^{3/2}N}e^{\alpha}\leq\frac{C}{N}\quad\text{when}\;\alpha>3/2.$ (3.48)
Now, Consider the fourth term of the R.H.S. of in the $y$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{c}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{c}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{y}\|_{\Omega_{c}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{c}}$
$\leq C\varepsilon^{1/2}\;h_{x}^{2}\|{F_{12}}_{yxx}\|_{\Omega_{c}}\;\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{5/2}}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;N$
$\leq C\frac{1}{\varepsilon^{2}N}e^{\alpha}\leq\frac{C}{N}\quad\text{when}\;\alpha>2.$ (3.49)
Applying similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})\leq C\frac{(\log(1/\varepsilon^{3/2})^{5/2}}{N^{3/2}}.$
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, we have
$\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
4 Estimates for convection and reaction terms involving discrete Green’s function
In this section, we aim to bound the convection and reaction portion of the bilinear form B(.,.) with its arguments parts of interpolation error uIuu-Iu and the discrete Green’s function $G_{mn}.$ Finally, based on the estimates for the diffusion, convection, and reaction terms, we derive the main theorem of this article.
Theorem 3
[8] Assume that $Iu$ represents the bilinear interpolation of $u$ on the Shishkin-type mesh. The interpolation error thus satisfies
$\|u-Iu\|_{L^{\infty}(\Omega_{f})}\leq CN^{-2}\log^{2}N,\quad\quad\|u-Iu\|_{L^{\infty}(\Omega_{c})}\leq CN^{-2}.$
Lemma 5.1 When $\beta\geq 2/3$ and the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Proof. We begin by decomposing the solution $u$ into its regular and singular parts as $u=R+F_{1}+F_{2}+F_{12}$. Consider the source point $(x_{m},y_{n})\in\Omega_{c}$. Hence,
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}=$ $(IR-R,b_{1}G_{mn,x})_{\Omega_{f,x}}+(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,x}}$
$+(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,x}}+(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,x}}$ (4.1)
Consider the first term of the R.H.S. of (4.1) $(IR-R,b_{1}G_{mn,x})_{\Omega_{f,x}}$.
$(IR-R,b_{1}G_{mn,x})_{\Omega_{f,x}}$ $\leq\|IR-R\|_{\Omega_{f,x}}\|b_{1}G_{mn,x}\|_{\Omega_{f,x}}$
$\leq C\|IR-R\|_{L^{\infty}(\Omega_{f,x})}\sqrt{|\Omega_{f,x}|}\;\|G_{mn,x}\|_{\Omega_{f,x}}$
$\leq\frac{C}{N^{2}}\sqrt{\varepsilon}\log^{1/2}(1/\varepsilon)\;\|G_{mn,x}\|_{\Omega_{f,x}}$
$\leq\frac{C}{N^{2}}\log^{1/2}(1/\varepsilon)\;\|G_{mn}\|_{1,\varepsilon}\leq\frac{C}{N}\log^{1/2}(1/\varepsilon).$ (4.2)
Consider the second term of the R.H.S. of (4.1) $(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,x}}$.
$(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,x}}$ $\leq\|IF_{1}-F_{1}\|_{\Omega_{f,x}}\|b_{1}G_{mn,x}\|_{\Omega_{f,x}}$
$\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{f,x})}\sqrt{|\Omega_{f,x}|}\;\|G_{mn,x}\|_{\Omega_{f,x}}$
$\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{f,x})}\sqrt{\varepsilon}\log^{1/2}(1/\varepsilon)\;\|G_{mn,x}\|_{\Omega_{f,x}}$
$\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{f,x})}\log^{1/2}(1/\varepsilon)\;\|G_{mn}\|_{1,\varepsilon}$
$\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{f,x})}\log^{1/2}(1/\varepsilon)\;\|N$
$\leq CN[h_{x}^{2}\|F_{1,xx}\|_{L^{\infty}(\Omega_{f,x})}+h_{x}h_{y}\|F_{1,xy}\|_{L^{\infty}(\Omega_{f,x})}+h_{y}^{2}\|F_{1,yy}\|_{L^{\infty}(\Omega_{f,x})}]$
$\leq CN\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}\frac{1}{\varepsilon^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon)}{N^{2}}\frac{1}{\varepsilon}+\frac{1}{N^{2}}\right]\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$ (4.3)
Next consider the third term of the R.H.S. of (4.1) $(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,x}}$.
$(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,x}}$ $\leq\|IF_{2}-F_{2}\|_{\Omega_{f,x}}\|b_{1}G_{mn,x}\|_{\Omega_{f,x}}$
$\leq C\|IF_{2}-F_{2}\|_{L^{\infty}(\Omega_{f,x})}\sqrt{|\Omega_{f,x}|}\;\|G_{mn,x}\|_{\Omega_{f,x}}$
$\leq C\|IF_{2}-F_{2}\|_{L^{\infty}(\Omega_{f,x})}\sqrt{\varepsilon}\log^{1/2}(1/\varepsilon)\;\|G_{mn,x}\|_{\Omega_{f,x}}$
$\leq C\|IF_{2}-F_{2}\|_{L^{\infty}(\Omega_{f,x})}\log^{1/2}(1/\varepsilon)\;\|G_{mn}\|_{1,\varepsilon}$
$\leq CN[h_{x}^{2}\|F_{2,xx}\|_{L^{\infty}(\Omega_{f,x})}+h_{x}h_{y}\|F_{2,xy}\|_{L^{\infty}(\Omega_{f,x})}+h_{y}^{2}\|F_{2,yy}\|_{L^{\infty}(\Omega_{f,x})}]$
$\leq CN\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon)}{N^{2}}\frac{1}{\varepsilon^{1/2}}\right]$
$\quad\quad\quad\quad\quad+CN\left[\frac{1}{N^{2}\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}\right].$ (4.4)
Consider
$\frac{1}{\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}\leq\frac{1}{\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}\|_{L^{\infty}(\Omega_{f,x})}+\frac{1}{\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}$ (4.5)
Note that $\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}\|_{L^{\infty}(\Omega_{f,x})}=\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}$.
Also let $\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}=e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y_{0})}$ where $y_{0}\in[-1+\lambda_{y},1-\lambda_{y}]$.
When $y_{0}\in[-1+\lambda_{y},1-\lambda_{y}]$, then
$y_{0}$ $\geq-1+\lambda_{y}\implies 1+y_{0}\geq\lambda_{y}$
$\implies\frac{-\beta}{\sqrt{\varepsilon}}(1+y_{0})$ $\leq\frac{-\beta}{\sqrt{\varepsilon}}\lambda_{y}\implies e^{\frac{-\beta}{\sqrt{\varepsilon}}(1+y_{0})}\leq e^{\frac{-\beta}{\sqrt{\varepsilon}}\lambda_{y}}\leq\varepsilon^{\frac{3\beta}{2}}.$ (4.6)
Similarly when $y_{0}\in[-1+\lambda_{y},1-\lambda_{y}]$, then
$y_{0}$ $\leq 1-\lambda_{y}\implies 1-y_{0}\geq\lambda_{y}$
$\implies\frac{-\beta}{\sqrt{\varepsilon}}(1-y_{0})$ $\leq\frac{-\beta}{\sqrt{\varepsilon}}\lambda_{y}\implies e^{\frac{-\beta}{\sqrt{\varepsilon}}(1-y_{0})}\leq e^{\frac{-\beta}{\sqrt{\varepsilon}}\lambda_{y}}\leq\varepsilon^{\frac{3\beta}{2}}.$ (4.7)
also for $x$, when $x\in[\lambda_{x},1]$
$x$ $\geq\lambda_{x}\implies\frac{-\alpha x}{\varepsilon}\leq\frac{-\alpha\lambda_{x}}{\varepsilon}$
$e^{\frac{-\alpha x}{\varepsilon}}$ $\leq e^{\frac{-\alpha\lambda_{x}}{\varepsilon}}=\varepsilon^{\alpha}.$ (4.8)
Combining (4.4 - 4.8), we obtain
$(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,x}}$ $\leq CN\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon)}{N^{2}}\frac{1}{\varepsilon^{1/2}}+\frac{\varepsilon^{\frac{3\beta}{2}-1}}{N^{2}}\right]$
$\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$ (4.9)
Next consider the fourth term of the R.H.S. of (4.1) $(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,x}}$.
$(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,x}}$ $\leq\|IF_{12}-F_{12}\|_{\Omega_{f,x}}\|b_{1}G_{mn,x}\|_{\Omega_{f,x}}$
$\leq C\|IF_{12}-F_{12}\|_{L^{\infty}(\Omega_{f,x})}\sqrt{|\Omega_{f,x}|}\;\|G_{mn,x}\|_{\Omega_{f,x}}$
$\leq C\|IF_{12}-F_{12}\|_{L^{\infty}(\Omega_{f,x})}\log^{1/2}(1/\varepsilon)\;\|G_{mn}\|_{1,\varepsilon}$
$\leq CN[h_{x}^{2}\|F_{12,xx}\|_{L^{\infty}(\Omega_{f,x})}+h_{x}h_{y}\|F_{12,xy}\|_{L^{\infty}(\Omega_{f,x})}$
$\quad\quad\quad\quad\quad+h_{y}^{2}\|F_{12,yy}\|_{L^{\infty}(\Omega_{f,x})}]$
$\leq\frac{C}{N}\left[\frac{\varepsilon^{2}\log^{2}(\frac{1}{\varepsilon})}{\varepsilon^{2}}+\|e^{\frac{-\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,x})}\frac{h_{x}}{\varepsilon^{3/2}}\right]$
$\quad\quad\quad\quad\quad+CN\left[\frac{1}{N^{2}\varepsilon}\|e^{\frac{-\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,x})}\right]$
$\leq C\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N}+\frac{\varepsilon\log(1/\varepsilon)}{N}\frac{\varepsilon^{\frac{3\beta}{2}}}{\varepsilon^{3/2}}+\frac{\varepsilon^{\frac{3\beta}{2}-1}}{N}\right]\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$ (4.10)
Hence when source point $(x_{m},y_{n})\in\Omega_{c}$ then we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Lemma 5.2 When $\alpha>2$ and the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$ and consider $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}$ but $u=R+F_{1}+F_{2}+F_{12}.$ Hence,
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}=$ $(IR-R,b_{1}G_{mn,x})_{\Omega_{f,y}}+(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,y}}$
$+$ $(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,y}}+(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,y}}$ (4.11)
Consider the second term of the R.H.S. of (4.11) $(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,y}}$.
$(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,y}}$ $\leq\|IF_{1}-F_{1}\|_{\Omega_{f,y}}\|b_{1}G_{mn,x}\|_{\Omega_{f,y}}$
$\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{f,y})}\sqrt{|\Omega_{f,y}|}\;\|G_{mn,x}\|_{\Omega_{f,y}}$
$\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{f,y})}\log^{1/2}(1/\varepsilon^{3/2})\varepsilon^{1/4}\;N\|G_{mn}\|_{\Omega_{f,y}}$
$\leq CN^{2}\varepsilon^{1/4}[h_{x}^{2}\|F_{1,xx}\|_{L^{\infty}(\Omega_{f,y})}+h_{x}h_{y}\|F_{1,xy}\|_{L^{\infty}(\Omega_{f,y})}$
$\quad+h_{y}^{2}\|F_{1,yy}\|_{L^{\infty}(\Omega_{f,y})}]$
$\leq CN^{2}\varepsilon^{1/4}\left[\frac{1}{N^{2}\varepsilon^{2}}+\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{\varepsilon N^{2}}+\varepsilon\log^{2}(1/\varepsilon^{3/2})\right]\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}$
$\leq\frac{C}{N}\quad\text{when}\;\alpha>2.$ (4.12)
Next consider the third term of the R.H.S. of (4.11) $(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,y}}$.
$(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,y}}$ $\leq\|IF_{2}-F_{2}\|_{\Omega_{f,y}}\|b_{1}G_{mn,x}\|_{\Omega_{f,y}}$
$\leq C\|IF_{2}-F_{2}\|_{L^{\infty}(\Omega_{f,y})}\sqrt{|\Omega_{f,y}|}\;\|G_{mn,x}\|_{\Omega_{f,y}}$
$\leq C\|IF_{2}-F_{2}\|_{L^{\infty}(\Omega_{f,y})}\log^{1/2}(1/\varepsilon^{3/2})\;\varepsilon^{1/4}N\|G_{mn}\|_{\Omega_{f,y}}$
$\leq CN^{2}\varepsilon^{1/4}[h_{x}^{2}\|F_{2,xx}\|_{L^{\infty}(\Omega_{f,y})}+h_{x}h_{y}\|F_{2,xy}\|_{L^{\infty}(\Omega_{f,y})}$
$\quad+h_{y}^{2}\|F_{2,yy}\|_{L^{\infty}(\Omega_{f,y})}]$
$\leq CN^{2}\varepsilon^{1/4}\left[\frac{1}{N^{2}}+\frac{\sqrt{}{\varepsilon}\log(1/\varepsilon^{3/2})}{N^{2}}\frac{1}{\varepsilon^{1/2}}+\frac{\varepsilon\log^{2}{(1/\varepsilon^{3/2})}}{\varepsilon N^{2}}\right]$
$\leq\frac{C\log^{2}({1/\varepsilon^{3/2}})}{N}.$ (4.13)
Next consider the fourth term of the R.H.S. of (4.11) $(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,y}}$.
$(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,y}}$ $\leq\|IF_{12}-F_{12}\|_{\Omega_{f,y}}\|b_{1}G_{mn,x}\|_{\Omega_{f,y}}$
$\leq C\|IF_{12}-F_{12}\|_{L^{\infty}(\Omega_{f,y})}\sqrt{|\Omega_{f,y}|}\;\|G_{mn,x}\|_{\Omega_{f,y}}$
$\leq C\|IF_{12}-F_{12}\|_{L^{\infty}(\Omega_{f,y})}\varepsilon^{1/4}\log^{1/2}(1/\varepsilon^{3/2})\;N\|G_{mn}\|_{\Omega_{f,y}}$
$\leq CN^{2}\varepsilon^{1/4}[h_{x}^{2}\|F_{12,xx}\|_{L^{\infty}(\Omega_{f,y})}+h_{x}h_{y}\|F_{12,xy}\|_{L^{\infty}(\Omega_{f,y})}$
$\quad\quad\quad\quad\quad+h_{y}^{2}\|F_{12,yy}\|_{L^{\infty}(\Omega_{f,y})}]$
$\leq CN^{2}\left[{\frac{\varepsilon^{1/4}}{\varepsilon^{2}N^{2}}}+\|e^{\frac{-\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,y})}\frac{h_{y}\varepsilon^{1/4}}{N^{2}\varepsilon^{3/2}}\right]$
$\;+CN^{2}\varepsilon^{1/4}\left[\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{N^{2}\varepsilon}\|e^{\frac{-\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,y})}\right]$
$\leq CN^{2}\varepsilon^{1/4}\left[\frac{1}{\varepsilon^{2}N^{2}}+\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\frac{\varepsilon^{\frac{3\beta}{2}}}{\varepsilon^{3/2}}+\frac{\varepsilon^{\frac{3\beta}{2}}}{N}\right]\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$ (4.14)
Consider the first term of the R.H.S. of (4.11) $(IR-R,b_{1}G_{mn,x})_{\Omega_{f,y}}$.
$(IR-R,b_{1}G_{mn,x})_{\Omega_{f,y}}$ $\leq\frac{C}{N}\quad\quad\quad\quad\quad\quad\quad\text{(By inverse inequality)}$ (4.15)
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}$ then $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}$
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Lemma 5.3 If the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}\leq C\frac{\log^{2}({1/\varepsilon^{3/2})}}{N}.$
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$ and consider $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}$ but $u=R+F_{1}+F_{2}+F_{12}.$ Hence,
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}=$ $(IR-R,b_{1}G_{mn,x})_{\Omega_{f,xy}}+(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,xy}}$
$+(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,xy}}+(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,xy}}$ (4.16)
Consider the first term of the R.H.S. of (4.16) $(IR-R,b_{1}G_{mn,x})_{\Omega_{f,xy}}$.
$(IR-R,b_{1}G_{mn,x})_{\Omega_{f,xy}}$ $\leq\|IR-R\|_{\Omega_{f,xy}}\|b_{1}G_{mn,x}\|_{\Omega_{f,xy}}$
$\leq C\|IR-R\|_{L^{\infty}(\Omega_{f,xy})}\sqrt{|\Omega_{f,xy}|}\;\|G_{mn,x}\|_{\Omega_{f,xy}}$
$\leq C\|IR-R\|_{L^{\infty}(\Omega_{f,xy})}\log(1/\varepsilon^{3/2})\varepsilon^{3/4}\;\|G_{mn}\|_{1,\varepsilon}$
$\leq\frac{CN}{\varepsilon^{-1/4}}[h_{x}^{2}\|R_{xx}\|_{L^{\infty}(\Omega_{f,xy})}+h_{x}h_{y}\|R_{xy}\|_{L^{\infty}(\Omega_{f,xy})}+h_{y}^{2}\|R_{yy}\|_{L^{\infty}(\Omega_{f,xy})}]$
$\leq\frac{CN}{\varepsilon^{-1/4}}\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}+\frac{\varepsilon\sqrt{\varepsilon}\log^{2}(1/\varepsilon^{3/2})}{N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{N^{2}}\right]$
$\leq C\frac{\log^{2}({1/\varepsilon^{3/2})}}{N}.$ (4.17)
Consider the second term of the R.H.S. of (4.16) $(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,xy}}$.
$(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,xy}}$ $\leq\|IF_{1}-F_{1}\|_{\Omega_{f,xy}}\|b_{1}G_{mn,x}\|_{\Omega_{f,xy}}$
$\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{f,xy})}\sqrt{|\Omega_{f,xy}|}\;\|G_{mn,x}\|_{\Omega_{f,xy}}$
$\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{f,xy})}\varepsilon^{3/4}\log(1/\varepsilon^{3/2})\;\|G_{mn,x}\|_{\Omega_{f,xy}}$
$\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{f,xy})}\varepsilon^{1/4}\log^{1/2}(1/\varepsilon)\;\|G_{mn}\|_{1,\varepsilon}$
$\leq CN[h_{x}^{2}\|F_{1,xx}\|_{L^{\infty}(\Omega_{f,xy})}+h_{x}h_{y}\|F_{1,xy}\|_{L^{\infty}(\Omega_{f,xy})}+h_{y}^{2}\|F_{1,yy}\|_{L^{\infty}(\Omega_{f,xy})}]$
$\leq CN\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}\varepsilon^{2}}+\frac{\varepsilon\sqrt{\varepsilon}\log^{2}(1/\varepsilon^{3/2})}{\varepsilon N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{N^{2}}\right]$
$\leq C\frac{\log^{2}({1/\varepsilon^{3/2})}}{N}.$ (4.18)
Next consider the third term of the R.H.S. of (4.16) $(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,xy}}$.
$(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,xy}}$ $\leq\|IF_{2}-F_{2}\|_{\Omega_{f,xy}}\|b_{1}G_{mn,x}\|_{\Omega_{f,xy}}$
$\leq C\|IF_{2}-F_{2}\|_{L^{\infty}(\Omega_{f,xy})}\sqrt{|\Omega_{f,xy}|}\;\|G_{mn,x}\|_{\Omega_{f,xy}}$
$\leq C\|IF_{2}-F_{2}\|_{L^{\infty}(\Omega_{f,xy})}\varepsilon^{1/4}\log(1/\varepsilon^{3/2})\;\|G_{mn}\|_{1,\varepsilon}$
$\leq CN[h_{x}^{2}\|F_{2,xx}\|_{L^{\infty}(\Omega_{f,xy})}+h_{x}h_{y}\|F_{2,xy}\|_{L^{\infty}(\Omega_{f,xy})}+h_{y}^{2}\|F_{2,yy}\|_{L^{\infty}(\Omega_{f,xy})}]$
$\leq CN\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}+\frac{\varepsilon\sqrt{\varepsilon}\log^{2}(1/\varepsilon^{3/2})}{\varepsilon^{1/2}N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{\varepsilon N^{2}}\right]$
$\leq C\frac{\log^{2}({1/\varepsilon^{3/2})}}{N}.$ (4.19)
Next consider the fourth term of the R.H.S. of (4.16) $(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,xy}}$.
$(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,xy}}$ $\leq\|IF_{12}-F_{12}\|_{\Omega_{f,xy}}\|b_{1}G_{mn,x}\|_{\Omega_{f,xy}}$
$\leq C\|IF_{12}-F_{12}\|_{L^{\infty}(\Omega_{f,xy})}\sqrt{|\Omega_{f,xy}|}\;\|G_{mn,x}\|_{\Omega_{f,xy}}$
$\leq C\|IF_{12}-F_{12}\|_{L^{\infty}(\Omega_{f,xy})}\varepsilon^{1/4}\log(1/\varepsilon^{3/2})\;\|G_{mn}\|_{1,\varepsilon}$
$\leq CN\varepsilon^{1/4}[h_{x}^{2}\|F_{12,xx}\|_{L^{\infty}(\Omega_{f,xy})}+h_{x}h_{y}\|F_{12,xy}\|_{L^{\infty}(\Omega_{f,xy})}$
$\quad\quad\quad\quad\quad+h_{y}^{2}\|F_{12,yy}\|_{L^{\infty}(\Omega_{f,xy})}]$
$\leq CN\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{\varepsilon^{2}N^{2}}+\frac{\varepsilon\sqrt{\varepsilon}\log^{2}(1/\varepsilon^{3/2})}{\varepsilon^{3/2}N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{\varepsilon N^{2}}\right]$
$\leq C\left[\frac{\log^{2}(1/\varepsilon)}{N}+\frac{\log^{2}(1/\varepsilon^{3/2})}{N}+\frac{\log^{2}(1/\varepsilon^{3/2})}{N}\right]$
$\leq C\frac{\log^{2}({1/\varepsilon^{3/2})}}{N}.$ (4.20)
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}$ then $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}$
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
Similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
Lemma 5.4 When $\alpha>5/2$ and the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $(Iu-u,b_{1}G_{mn,x})_{\Omega_{c}}\leq C\frac{\log^{2}({1/\varepsilon^{3/2})}}{N}.$
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$ and consider $(b_{1}(Iu-u)_{x},G_{mn})_{\Omega_{c}}$.
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{c}}=$ $(IR-R,b_{1}G_{mn,x})_{\Omega_{c}}+(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{c}}$
$+(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{c}}+(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{c}}$ (4.21)
Consider the first term of the R.H.S. of (4.21) $(IR-R,b_{1}G_{mn,x})_{\Omega_{c}}$.
$(IR-R,b_{1}G_{mn,x})_{\Omega_{c}}$ $\leq\|IR-R\|_{\Omega_{c}}\|b_{1}G_{mn,x}\|_{\Omega_{c}}\leq C\|IR-R\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;\|G_{mn,x}\|_{\Omega_{c}}$
$\leq C\|IR-R\|_{L^{\infty}(\Omega_{c})}\;\varepsilon^{-1/2}\|G_{mn}\|_{1,\varepsilon}$
$\leq\frac{CN}{\sqrt{\varepsilon}}[h_{x}^{2}\|R_{xx}\|_{L^{\infty}(\Omega_{c})}+h_{x}h_{y}\|R_{xy}\|_{L^{\infty}(\Omega_{c})}+h_{y}^{2}\|R_{yy}\|_{L^{\infty}(\Omega_{c})}]$
$\leq\frac{CN}{\sqrt{\varepsilon}}\left[\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{N^{2}}\right]$
$\leq\frac{C}{N}\quad\text{when}\quad\alpha>1/2.$ (4.22)
Consider the second term of the R.H.S. of (4.21) $(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{c}}$.
$(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{c}}$ $\leq\|IF_{1}-F_{1}\|_{\Omega_{c}}\|b_{1}G_{mn,x}\|_{\Omega_{c}}\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;\|G_{mn,x}\|_{\Omega_{c}}$
$\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{c})}\;\|G_{mn,x}\|_{\Omega_{c}}$
$\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{c})})\;\varepsilon^{-1/2}\|G_{mn}\|_{1,\varepsilon}$
$\leq\frac{CN}{\sqrt{\varepsilon}}[h_{x}^{2}\|F_{1,xx}\|_{L^{\infty}(\Omega_{c})}+h_{x}h_{y}\|F_{1,xy}\|_{L^{\infty}(\Omega_{c})}+h_{y}^{2}\|F_{1,yy}\|_{L^{\infty}(\Omega_{c})}]$
$\leq\frac{CN}{\sqrt{\varepsilon}}\left[\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{\varepsilon^{2}N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{\varepsilon N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{N^{2}}\right]$
$\leq\frac{C}{N}\quad\text{when}\;\alpha>5/2.$ (4.23)
Next consider the third term of the R.H.S. of (4.21) $(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{c}}$.
$(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{c}}$ $\leq\|IF_{2}-F_{2}\|_{\Omega_{c}}\|b_{1}G_{mn,x}\|_{\Omega_{c}}$
$\leq C\|IF_{2}-F_{2}\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;\|G_{mn,x}\|_{\Omega_{c}}$
$\leq C\|IF_{2}-F_{2}\|_{L^{\infty}(\Omega_{c})}\;\varepsilon^{-1/2}\|G_{mn}\|_{1,\varepsilon}$
$\leq\frac{CN}{\sqrt{\varepsilon}}[h_{x}^{2}\|F_{2,xx}\|_{L^{\infty}(\Omega_{c})}+h_{x}h_{y}\|F_{2,xy}\|_{L^{\infty}(\Omega_{c})}+h_{y}^{2}\|F_{2,yy}\|_{L^{\infty}(\Omega_{c})}]$
$\leq\frac{CN}{\sqrt{\varepsilon}}\left[\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{\varepsilon^{1/2}N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{\varepsilon N^{2}}\right]$
$\leq\frac{C}{N}\quad\text{when}\;\alpha>3/2.$ (4.24)
Next consider the fourth term of the R.H.S. of (4.21) $(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{c}}$.
$(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{c}}$ $\leq\|IF_{12}-F_{12}\|_{\Omega_{c}}\|b_{1}G_{mn,x}\|_{\Omega_{c}}$
$\leq C\|IF_{12}-F_{12}\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;\|G_{mn,x}\|_{\Omega_{c}}$
$\leq C\|IF_{12}-F_{12}\|_{L^{\infty}(\Omega_{c})}\varepsilon^{-1/2}\;\|G_{mn}\|_{1,\varepsilon}$
$\leq CN\varepsilon^{-1/2}[h_{x}^{2}\|F_{12,xx}\|_{L^{\infty}(\Omega_{c})}+h_{x}h_{y}\|F_{12,xy}\|_{L^{\infty}(\Omega_{c})}$
$\quad+h_{y}^{2}\|F_{12,yy}\|_{L^{\infty}(\Omega_{c})}]\leq\frac{CN}{\varepsilon^{1/2}}\left[\frac{\varepsilon^{\alpha}}{\varepsilon^{2}N^{2}}+\frac{\varepsilon^{\alpha}}{\varepsilon^{3/2}N^{2}}+\frac{\varepsilon^{\alpha}}{\varepsilon N^{2}}\right]$
$\leq C\left[\frac{\varepsilon^{\alpha-\frac{5}{2}}}{N}\right]\leq\frac{C}{N}\quad\text{when}\;\alpha>5/2.$ (4.25)
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}$ then we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{c}}\leq\frac{C}{N}.$
Similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{c}}\leq\frac{C}{N}.$
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{c}}\leq\frac{C}{N}.$
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
Similarly arguments yields that when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then we have
$(Iu-u,b_{1,x}G_{mn})_{\Omega}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
Lemma 5.5 If the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $(c(Iu-u),G_{mn}).\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
[…]
$(c(Iu-u),G_{mn})_{\Omega}$ $=(c(Iu-u),G_{mn})_{\Omega_{c}}$
$+(c(Iu-u),G_{mn})_{\Omega_{f,x}}$
$+(c(Iu-u),G_{mn})_{\Omega_{f,y}}$
$+(c(Iu-u),G_{mn})_{\Omega_{f,xy}}$ (4.26)
Similar arguments yields that when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$ then the reaction term bounds by
$(c(Iu-u),G_{mn})_{\Omega}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
Theorem 4
When $\alpha\geq 5/2$ and $\beta\geq 2/3$ and $u_{h}$ denotes the approximate solution of the problem, and the source point $(x_{m},y_{n})$ only lies in the coarse region $\Omega_{c}$ and $x$-layer region $\Omega_{f,x}$. Building upon the preceding analysis, the following estimate holds
$|(u-u_{h})(x_{m},y_{n})|\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$
Proof. At the beginning, we aim to consider $|(Iu-u_{h})(x_{m},y_{n})|$
Let $(x_{m},y_{n})$ be the mesh point then
$(Iu-u_{h})(x_{m},y_{n})=(\delta_{mn},Iu-u_{h})$ (4.27)
Where $\delta_{mn}$ is the Kronecker delta function defined on mesh point $(x_{m},y_{n})$. Let $G_{mn}$ be the discrete Green’s function corresponding to the differential operator coming from problem (1.2) and the discrete Dirac-Delta function $\delta_{mn}$, then
$(Iu-u_{h})(x_{m},y_{n})=(\delta_{mn},Iu-u_{h})=B(Iu-u_{h},G_{mn})$ (4.28)
where
$B(u,v)=\varepsilon\int_{\Omega}\nabla u.\nabla v\;dx+\int_{\Omega}b_{1}u_{x}v\;dx+\int_{\Omega}cuv\;dx$
$B(u,v)=\varepsilon(\nabla u,\nabla v)+(b_{1}u_{x},v)+(cu,v)$ (4.29)
$\implies(Iu-u_{h})(x_{m},y_{n})=B(Iu-u,G_{mn})+B(u-u_{h},G_{mn})$ (4.30)
But by the Galerkin’s orthogonality, we can write that
$B(u-u_{h},G_{mn})=0$ (4.31)
$\implies(Iu-u_{h})(x_{m},y_{n})=B(Iu-u,G_{mn})$ (4.32)
$\implies(Iu-u_{h})(x_{m},y_{n})=\varepsilon\int_{\Omega}\nabla(Iu-u).\nabla G_{mn}\;dxdy\\ +\int_{\Omega}b_{1}(Iu-u)_{x}G_{mn}\;dxdy+\int_{\Omega}c(Iu-u)G_{mn}\;dxdy$
$\implies(Iu-u_{h})(x_{m},y_{n})=\varepsilon(\nabla(Iu-u),\nabla G_{mn})+(b_{1}(Iu-u)_{x},G_{mn})\\ +(c(Iu-u),G_{mn})$ (4.33)
In this theorem, we aim to bound the bilinear form B(.,.) appearing in (4.29) with its arguments the interpolation error uIuu-Iu and the discrete Green’s function $G_{mn}.$ We decompose the solution $u$ into regular and singular parts, where we note that the singular part denotes the part of the solution that contains layers. Therefore, the singular part can be decomposed further into three more regions, namely $F_{1}$, $F_{2}$, and $F_{12}$, where $F_{1}$ contains the interior layer, $F_{2}$ contains the boundary layer, and $F_{12}$ contains both the interior and boundary layers in the corner region.
To establish error estimates, we have to estimate the term on the R.H.S. of (4.33), we begin in the first term of the R.H.S. of (4.33) is $\varepsilon(\nabla(Iu-u),\nabla G_{mn})$ and decompose the solution $u=R+F_{1}+F_{2}+F_{12}$, we obtain
$\varepsilon(\nabla(u-Iu),\nabla G_{mn})=\varepsilon(\nabla((R+F_{1}+F_{2}+F_{12})-I(R+F_{1}+F_{2}+F_{12})),G_{mn})\\ =\varepsilon(\nabla(R-IR),\nabla G_{mn})+\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})\\ +\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})+\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})$ (4.34)
Using Lemma (4.1 to 4.4), we have that when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then the diffusion term is bounded by
$(\nabla(u-Iu),\nabla G_{mn})_{\Omega}\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$ (4.35)
Now, the second term of the R.H.S of (4.33) is $(b_{1}(Iu-u)_{x},G_{mn})$, using integration by parts, we get
$(b_{1}(Iu-u)_{x},G_{mn})=$ $(b_{1}(Iu-u),G_{mn})_{\partial\Omega}-(Iu-u,b_{1}G_{mn,x})_{\Omega}$
$(b_{1}(Iu-u)_{x},G_{mn})=$ $-(Iu-u,b_{1}G_{mn,x})_{\Omega}.$
and using Lemma (5.1 to 5.4), we have that when the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$ then the convection term is bounded by
$(b_{1}(Iu-u)_{x},G_{mn})_{\Omega}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$ (4.36)
Now, the third term of the R.H.S. of (4.33) is $(c(Iu-u),G_{mn})$ and using Lemma 5.5 we have when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$ the reaction term bounds by
$(c(Iu-u),G_{mn})_{\Omega}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$ (4.37)
Hence, combining (4.35), (4.36), and (4.37) completes the proof.
5 Numerical results and discussion
In this section, we take a numerical example to validate the theoretical results established in Theorems (4). We apply our method to a test problem to verify the uniform rate of convergence only in the $x$-layer and coarse layer regions.
5.1 Numerical implementation
We evaluate the error and the uniform rate of convergence with respect to the maximum norm in the coarse layer and $x$-layer regions for the different values of the perturbation parameters $\varepsilon$ and the discretization parameter N. In the absence of an exact solution to the text example, we employ the double mesh principle technique to compute the error and order of the convergence. We calculate the error estimates in the maximum norm using $\|U_{N}-U_{2N}\|_{L^{\infty}{(\Omega})}$. The convergence rate is calculated by using the following:
$C^{N}_{\varepsilon}=\log_{2}\left(\frac{\|U_{N}-U_{2N}\|_{L^{\infty}(\Omega)}}{\|U_{2N}-U_{4N}\|_{L^{\infty}(\Omega)}}\right)$
A numerical example is presented where the exact solution is not available.
Example 5.1 We investigate the turning point singularly perturbed equation provided below
$\begin{split}-\varepsilon\Delta u-x(x^{2}+e^{1+xy})u_{x}+(3+x^{2}e^{x})u&=f(x,y),\quad\text{in}\;\Omega=(-1,1)\times(-1,1),\\ u&=0,\quad\Gamma=\partial\Omega\end{split}$ (5.1)
For this particular problem, the source term is set as $f(x,y)=\frac{xy}{1+x^{2}+y^{2}}$.
A key feature of this problem is its turning point at $x=0$, which gives rise to an interior layer along the $x$-axis. In parallel, two boundary layers are present at the bottom of the boundary y=1y=-1 and the top of the boundary $y=1$ of the domain. A double-mesh principle is used to improve the accuracy and convergence behavior of the numerical scheme. This scheme allows us to estimate errors and evaluate the rate of convergence for different values of the perturbation parameter $\varepsilon$ and the mesh discretization parameter N without knowing the exact solution. The resultant error values and computed rates of convergence are systematically presented in Tables.
Table 1: Error in the subdomain $\Omega_{c}$ for example (5.1).
Error in $\Omega_{c}$
Number of discretization parameter NN
$\varepsilon$ 16 32 64 128 256
$10^{-5}$ 5.035E-02 3.512E-02 2.167E-02 1.270E-02 6.766E-03
$10^{-6}$ 5.315E-02 3.881E-02 2.463E-02 1.469E-02 8.779E-03
$10^{-7}$ 5.294E-02 4.033E-02 2.639E-02 1.597E-02 9.604E-03
$10^{-8}$ 5.173E-02 4.088E-02 2.757E-02 1.695E-02 1.022E-02
$10^{-9}$ 5.027E-02 4.100E-02 2.844E-02 1.777E-02 1.075E-02
Table 2: Error in the subdomain $\Omega_{f,x}$ for example (5.1).
Error in $\Omega_{f,x}$
Number of discretization parameter NN
$\varepsilon$ 16 32 64 128 256
$10^{-5}$ 1.008E-01 5.169E-02 2.684E-02 1.445E-02 7.501E-03
$10^{-6}$ 1.049E-01 5.786E-02 2.875E-02 1.597E-02 9.269E-03
$10^{-7}$ 1.068E-01 6.295E-02 2.978E-02 1.709E-02 9.969E-03
$10^{-8}$ 1.075E-01 6.714E-02 3.158E-02 1.812E-02 1.057E-02
$10^{-9}$ 1.077E-01 7.060E-02 3.449E-02 1.952E-02 1.114E-02
[…]
Table 3: Rate of convergence in the subdomains $\Omega_{c}$ and $\Omega_{f,x}$ for test example (5.1).
$\Omega_{c}$ $\Omega_{f,x}$
Number of discretization parameter NN Number of discretization parameter NN
$\varepsilon$ 16 32 64 128 16 32 64 128
$10^{-5}$ 0.5198 0.6965 0.7710 0.9083 0.9635 0.9452 0.8938 0.9458
$10^{-6}$ 0.4535 0.6560 0.7456 0.7429 0.8591 1.0089 0.8484 0.7847
$10^{-7}$ 0.3926 0.6116 0.7246 0.7339 0.7627 1.0799 0.8015 0.7773
$10^{-8}$ 0.3396 0.5683 0.7020 0.7303 0.6795 1.0884 0.8011 0.7778
$10^{-9}$ 0.2941 0.5277 0.6788 0.7254 0.6093 1.0334 0.8214 0.8094
Table 4: Error in the subdomain $\Omega_{f,y}$ for example (5.1).
Error in $\Omega_{f,y}$
Number of discretization parameter NN
$\varepsilon$ 16 32 64 128 256
$10^{-5}$ 9.021E-02 9.552E-02 9.105E-02 8.140E-02 6.415E-02
$10^{-6}$ 8.718E-02 9.527E-02 9.300E-02 8.630E-02 7.804E-02
$10^{-7}$ 8.173E-02 9.177E-02 9.157E-02 8.630E-02 7.956E-02
$10^{-8}$ 7.624E-02 8.753E-02 8.918E-02 8.526E-02 7.932E-02
$10^{-9}$ 7.133E-02 8.340E-02 8.659E-02 8.395E-02 7.874E-02
Table 5: Error in the subdomain $\Omega_{f,xy}$ for example (5.1).
Error in $\Omega_{f,xy}$
Number of discretization parameter NN
$\varepsilon$ 16 32 64 128 256
$10^{-5}$ 1.600E-01 1.506E-01 1.040E-01 8.613E-02 6.732E-02
$10^{-6}$ 1.514E-01 1.525E-01 9.997E-02 8.824E-02 7.935E-02
$10^{-7}$ 1.432E-01 1.516E-01 1.066E-01 8.797E-02 8.004E-02
$10^{-8}$ 1.360E-01 1.496E-01 1.130E-01 8.797E-02 7.949E-02
$10^{-9}$ 1.298E-01 1.470E-01 1.179E-01 8.921E-02 7.880E-02
Table 6: Rate of convergence in the subdomains $\Omega_{y}$ and $\Omega_{f,xy}$ for test example (5.1).
$\Omega_{y}$ $\Omega_{f,xy}$
Number of discretization parameter NN Number of discretization parameter NN
$\varepsilon$ 16 32 64 128 16 32 64 128
$10^{-5}$ -0.0825 0.0691 0.1617 0.3435 0.0873 0.5344 0.2721 0.3556
$10^{-6}$ -0.1280 0.0347 0.1079 0.1450 -0.0098 0.6088 0.1800 0.1533
$10^{-7}$ -0.1672 0.0032 0.0854 0.1174 -0.0821 0.5078 0.2772 0.1364
$10^{-8}$ -0.1994 -0.0268 0.0648 0.1043 -0.1371 0.4044 0.3612 0.1462
$10^{-9}$ -0.2254 -0.0543 0.0448 0.0924 -0.1798 0.3190 0.4016 0.1790
5.2 Validation
For text example (5.1), Tables (1), (2), (4) and (5) display the numerical errors for subdomains $\Omega_{c}$, $\Omega_{f,x}$, $\Omega_{f,y}$ and $\Omega_{f,xy}$ respectively. Additionally, Tables (3) and (6) present the rates of convergence when the source point is located within $\Omega_{c}\cup\Omega_{f,x}$ and $\Omega_{f,y}\cup\Omega_{f,xy}$, respectively. As shown in Table (3), the proposed scheme is uniformly convergent for the source point that lies in $\Omega_{c}\cup\Omega_{f,x}$. Conversely, Table (6) indicates that uniformity is lost when the source point lies in $\Omega_{f,y}\cup\Omega_{f,xy}$. The solution plot clearly captures the interior and boundary layers. Furthermore, the numerical convergence rates, which are linear up to a logarithmic factor when the source point lies within the subdomain $\Omega_{c}\cup\Omega_{f,x}$, confirm our theoretical predictions.
6 Conclusions
In this article, we employ a standard finite element method to solve a two-dimensional singularly perturbed convection-diffusion turning point problem. Because the convection coefficient change sign within the domain, the problem has a interior layer along $x$-axis and two boundary layers along $y$-axis. Parameter uniform error estimates are obtained for the coarse and $x$ layer regions using Shishkin-type mesh with specific mesh transition parameter. Additionally, pointwise error estimates are established through the application of the discrete Green’s function. Stynes analysis indicated, uniformly convergent in the perturbation parameter of order $N^{-1}\log N$ in global energy norm and of order $N^{-1/2}\log^{3/2}N$ pointwise near the outflow boundary but not in the coarse region. We establish uniform convergence in the perturbation parameter of order $N^{-1}\log^{3}(1/\varepsilon^{3/2})$ pointwise in the coarse and $x$-layer regions.
Acknowledgments
The authors are thankful to the anonymous referees for their valuable comments and suggestions to improve the paper. The authors thank LNMIIT, Jaipur for providing the facilities and support necessary to complete this research. The authors also acknowledge the DST-FIST program (Govt. of India) for providing the financial support for setting up the computing lab facility under the scheme ”Fund for Improvement of Science and Technology” (FIST - No. SR/EST/MS-1/2018/24). The authors also acknowledge the technical support provided by the LNM Institute of Information Technology, India.